OpenStudy (anonymous):

What is the radical exponent of:

OpenStudy (anonymous):

\[^3\sqrt{x^3}\]

OpenStudy (anonymous):

Really need the answer ASAP. Thanks so much! @smileyxl3 @sara17 @AravindG @funinabox

OpenStudy (jdoe0001):

what do they mean by " radical exponent of" ?

OpenStudy (anonymous):

you can write it like this: (x^3)^(1/3) in generality: \[\sqrt[n]{x}\] is writen as: (x)^(1/n) then you should multiply powers.. so: (x)^(3)(1/3)=x because (3)(1/3)=1 so => x^1=x

OpenStudy (anonymous):

I'm sorry I meant simplify it

OpenStudy (anonymous):

Thanks so much @sara17 Can you do another one please? Simplify: \[\sqrt[11]{x^5 * x^6}\]

OpenStudy (jdoe0001):

\(\huge \sqrt[3]{x^3} \) if the exponent inside the radical, MATCHES the root of the radical, then it comes out, without an exponent, just by itself :)

OpenStudy (jdoe0001):

what's \(\bf \large x^5 \times x^6 \quad ?\)

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg

OpenStudy (anonymous):

just multiply powers so you will have: (x^5)(x^6)=(x)^(5+6)=x^11 then you will have: \[\sqrt[11]{x^11}\] then write it in this form: (x^11)^(1/11)=x^((11)(1/11))=x^1=x

OpenStudy (anonymous):

sorry i wrote multiply but i meant sum of powers: 5+6=11